OpenStudy (anonymous):

70 = –7(–2 – 2z) z = 4 z = –28 z = –112 z = 784

OpenStudy (anonymous):

70/-7=-2-2z -10=-2-2z -10+2=-2z -8=-2z z=-8/-2

OpenStudy (anonymous):

z=4

OpenStudy (anonymous):

thannks